Lauren knows she can afford to make monthly payments of $575 for 36 months. How much will the bank lend her today on a 3% APR car loan in exchange for her promised monthly payments?
Answer #1

Ans $ 19772.22

P = Regular Payments
PV = Loan Amount
r = rate of interest
n = no of periods
P = r (PV)
1 - (1 + r )-n
575 = (3%/12)*PV
1 - (1 / (1 + 3%/12)^36))
575 = 0.0025 * PV
0.085966163
PV = 575 * 0.085966163 / 0.0025
PV = 19772.22